OpenStudy (anonymous):

using Laplace: x''+4x'+5x=e^t , x(0)=1 , x'(0)=2 I'm literally just stuck entirely on this problem, and can't find enough resources to work through it.

OpenStudy (anonymous):

Lapalace method?!

OpenStudy (anonymous):

LaPlace Transform Method

OpenStudy (amistre64):

a laplace table would be useful

OpenStudy (amistre64):

\[L\{ x''+4x'+5x=e^t\}\] \[L\{ x''\}+4L\{x'\}+5L\{x\}=L\{e^t\}\]is the easy setup at least :)

OpenStudy (anonymous):

I have simplified it down to the inverse Laplace form (I believe) \[(e ^{t} + S + 6)/((S+2)^2-9)\] Then where to?

OpenStudy (amistre64):

you cant have a "t" in the outcome there

OpenStudy (amistre64):

and if you ike prettier fractions type in frac{a}{b} in teh equation editor; or the latex markups in the reply box

OpenStudy (amistre64):

no

OpenStudy (anonymous):

Oh, yea. I just noticed that as well. Do I treat the =e^t as the function, and set it equal to zero?

OpenStudy (amistre64):

no

OpenStudy (anonymous):

1/s-1

OpenStudy (amistre64):

treat it as part of the whole shindig

OpenStudy (amistre64):

lets go thru the long versions for practice, cause i need the practice lol

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Firing up LaTeX so I can do Equations

OpenStudy (amistre64):

L{x''} = :e^-st x'' = x' e^-st +s :e^-st x' = x' e^-st +s (e^-st x + s: e^-st x) = x' e^-st +s (e^-st x + sX(s)) = x' e^-st +se^-st x + s^2X(s) ; (0 , inf) = s^2X(s) -x'(0) -sx(0) right?

OpenStudy (anonymous):

you may be coming at it from a different angle than me. I could always be wrong though. I ended up with: \[L{x}(S ^{2}+4S+5) -S -6 = e^t\]

OpenStudy (amistre64):

\[s^2X(s)-s-2+4sX(s)-4+5X(s)=\frac{1}{s-1}\] \[X(s)(s^2+4s+5)=\frac{1}{s-1}+6+s\] \[X(s)=\frac{\frac{1}{s-1}+6+s}{s^2+4s+5}\]

OpenStudy (anonymous):

that looks 100 times better!

OpenStudy (amistre64):

i gotta chk my e^t tho

OpenStudy (amistre64):

s+1 not s-1 lol

OpenStudy (amistre64):

simplify that mess, and then we can work on undoing it to get a solution

OpenStudy (anonymous):

Then I did complete the square to find the bottom.

OpenStudy (anonymous):

So right now I have your numerator over (S+2)^2 - 9

OpenStudy (amistre64):

\[X(s)=\frac{1+6(s+1)+s(s+1)}{(s+1)(s^2+4s+5)}\] \[X(s)=\frac{1+6s+6+s^2+s}{(s+1)(s^2+4s+5)}\] \[X(s)=\frac{s^2+7s+7}{(s+1)(s^2+4s+5)}\]

OpenStudy (anonymous):

Yup, that's where I am now. Time for conversion you think?

OpenStudy (amistre64):

not yet, we have to split this up into partial fractions in order to better format this into something invertible

OpenStudy (anonymous):

I don't think you can.

OpenStudy (anonymous):

nvm.

OpenStudy (amistre64):

\[\frac{s^2+7s+7}{(s+1)(s^2+4s+5)}=\frac{a}{(s+1)}+\frac{bx+c}{(s^2+4s+5)}\] \[\frac{-1^2+7.-1+7}{\cancel{(s+1)}(-1^2+4.-1+5)}=\frac{a}{(s+1)}+\frac{bx+c}{(s^2+4s+5)}\] \[\frac{1-7+7}{1-4+5}=\frac{a=1/2}{(s+1)}+\frac{bx+c}{(s^2+4s+5)}\]

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

Wait, you got A as 1/2? I have it as -2.

OpenStudy (amistre64):

double chk your math

OpenStudy (amistre64):

im papering it at the moment

OpenStudy (amistre64):

i get 4s+18 for the bs+c parts

OpenStudy (anonymous):

I hit complex roots -.-

OpenStudy (amistre64):

complex roots are fine

OpenStudy (anonymous):

I know, but I have had them on every question so far.

OpenStudy (anonymous):

hmm, I think i've ended up with: 1/8 e^(-5t)(-e^-4t)+10e^6t-1)

OpenStudy (anonymous):

\[\frac18 e ^{-5t}(-e ^{4t}+10e ^{6t}-1)\]

OpenStudy (amistre64):

you think thats good?

OpenStudy (anonymous):

It looks right, do you have anything?

OpenStudy (anonymous):

Plus, I need to eat dinner and the dining hall closes very soon.

OpenStudy (amistre64):

yeah,check this out http://www.stanford.edu/~boyd/ee102/laplace-table.pdf i think 1/(s-1) was correct to begin with so these might be bad if we used my stuff

OpenStudy (amistre64):

the rest was fine tho lol

OpenStudy (anonymous):

alright. Well, thanks a million for your help.